Assuming that Coca Cola has the same specific heat as water ( 4.18 J/(g x degrees C) ), calculate the amount of heat in J transferred when one can ( about 250g) is cooled from 15 degrees C to 6 degrees C . *
1
Expert's answer
2020-09-05T14:41:10-0400

"\u2206H = mC\u2206T"

"m = 250 g"

"C = 4.18 \frac{J}{(g\times°C)}"

"\u2206T = 15 - 6 = 9 \u00b0C"

"∆H = 250 g \times 4.18 \frac{J}{(g\times°C)} \times 9 °C = 9405 J"

Answer: 9405 J